# Sequential Monte Carlo Squared for Online Inference in Stochastic Epidemic Models

### v1

@@ Appendix B Posterior estimate in the simulation study
Table B.2: Summary of parameters in Experiment 2. Posterior median and 95% credible intervals using O-SMC2 with varying window sizes $t_{k}$ at the final time step $T=100$. “Relative Efficiency ” denotes the computational time reduction factor compared to SMC2.
@@
While smaller $t_{k}$ values yield significant computational savings, this efficiency advantage decreases as $t_{k}$ increases. For most tested window sizes, the credible intervals align closely with those from SMC2, indicating minimal bias in this application. However, smaller windows may introduce minor biases if the window does not capture sufficient temporal dependencies, potentially resulting in slightly larger 95% credible intervals.
Figure B.1: Parameters’ posterior densities in Experiment 1 corresponding to different window sizes $t_{k}$.
Figure B.2: Parameters’ posterior densities in Experiment 2 corresponding to different window sizes $t_{k}$
It is important to note that the windowing mechanism is applied only during the resampling step (i.e., when ESS drops below 50% of the total number of parameter particles in our case). Thus, the greater the frequency of resampling at time $t$ (especially when $t$ is significantly larger than $t_{k}$), the more computationally beneficial O-SMC2 becomes. In Figure B.3, we observe that the frequency of resampling at times well beyond $t_{k}=20$ gradually decreases in Experiment 1, while it remains more frequent in Experiment 2. This explains the greater computational gain observed in Experiment 2.
Figure B.3: Effective Sample Size. The left graph shows the the evolution of the ESS in Experiment 1 and the right one the evolution in Experiment 2.
Appendix C Parameter settings in the COVID-19 model
Table C.1 summarizes parameters from the COVID-19 model (Equation (3.2.1)), including hyperparameters (Equations (20) and (23)). For each parameter, prior and posterior estimates (median and 95% credible interval) are provided. Priors were informed by COVID-19 dynamics literature in Ireland (Irish Department of Health, 2020; Cazelles et al., 2021).
[…]
@@ Appendix C Parameter settings in the COVID-19 model
$\beta_{0}$ Initial condition transmission rate $\mathcal{U}(0.6,0.8)$
@@
Figure C.1: Daily estimates of the time-dependent transmission rate and effective reproduction number. The solid black lines represent the median estimate of $R_{\text{eff}}(t)$, with shaded blue regions indicating posterior uncertainty quantiles. Weekly averaged estimates of $R_{\text{eff}}(t)$ are shown in Figure 5.
Appendix D Application in real-time surveillance
The methodology developed in this paper is applied to a real-time surveillance setting, demonstrating how our model can track evolving dynamics in sequential data. By integrating new observations over time, the model provides insights into the system’s behavior. The focus is on an adaptive framework that updates the posterior distribution, conditioned on incoming data, while utilizing results from previous O-SMC2 runs as prior information. The results are shown in Figure D.1, where we observe that there is no significant difference compared to the results obtained from running the model from scratch (see Figure 5).
Figure D.1: Real-time application of O-SMC2 to sequential data: (Row 1) February 29, 2020, to May 30, 2020; (Row 2) May 30, 2020, to August 29, 2020; (Row 3) August 29, 2020, to November 28, 2020; and (Row 4) November 28, 2020, to February 27, 2021. Each row depicts the sequential estimates of daily new cases and the weekly average effective reproduction number ($R_{\text{eff}}(t)$). The solid black lines represent the median estimate, while the shaded blue regions indicate posterior uncertainty quantiles.

### v2

@@ Appendix C Posterior estimate in the simulation study
Table 3: Summary of parameter estimates in Experiment 2. Posterior medians and 95% credible intervals were obtained using O-SMC2 with varying window sizes $t_{k}$ at the final time step $T=100$. “Relative Efficiency” denotes the factor by which computational time is reduced compared to SMC2.
@@
While smaller $t_{k}$ values yield significant computational savings, this efficiency advantage decreases as $t_{k}$ increases. For most tested window sizes, the credible intervals align closely with those from SMC2, indicating minimal bias in this application. However, smaller windows may introduce minor biases if the window does not capture sufficient temporal dependencies, potentially resulting in slightly larger 95% credible intervals (Figures 1-2).
Figure 1: Posterior densities of parameters at $T=100$ (top) and filtered estimates (bottom) in Experiment 1. The dashed lines in the bottom plot indicate the 95% credible intervals. True parameter values are represented by orange dashed lines.
Figure 2: Posterior densities of parameters at $T=100$ (top) and filtered estimates (bottom) in Experiment 2. The dashed lines in the bottom plot indicate the 95% credible intervals. True parameter values are represented by orange dashed lines.
It is important to note that the windowing mechanism is applied only during the resampling step (i.e., when ESS drops below 50% of the total number of parameter particles in our case). Thus, the greater the frequency of resampling at time $t$ (especially when $t$ is significantly larger than $t_{k}$), the more computationally beneficial O-SMC2 becomes. In Figure 3, we observe that the frequency of resampling at times well beyond $t_{k}=20$ gradually decreases in Experiment 1, while it remains more frequent in Experiment 2. This explains the greater computational gain observed in Experiment 2.
Figure 3: Effective Sample Size. The left graph shows the the evolution of the ESS in Experiment 1 and the right one the evolution in Experiment 2.
Appendix D Parameter settings in the COVID-19 model
Table 1 summarizes parameters from the COVID-19 model (Equation (3.2.1)), including hyperparameters (Equations (20) and (23)). For each parameter, prior and posterior estimates (median and 95% credible interval) are provided. Priors were informed by COVID-19 dynamics literature in Ireland (Irish Department of Health, 2020; Cazelles et al., 2021).
Table 1: Description of the different parameters, priors and posteriors estimate at the final time $T=365$ days. Upper bound and/or lower bound have been imposed by the observations. $\mathcal{TN}$ stands for truncated normal distribution ($\mathcal{TN}_{[\text{inf, sup}]}(\text{mean, std})$).
@@ Appendix D Parameter settings in the COVID-19 model
$\beta(0)$ Initial condition transmission rate $\mathcal{U}(0.6,0.8)$ -
@@
To evaluate predictive accuracy, we report the Mean Absolute Error (MAE) and the Continuous Ranked Probability Score (CRPS).
The MAE quantifies the average absolute deviation between the observed values and the model’s predictive mean:
$\text{MAE}=\frac{1}{T}\sum_{t=1}^{T}|y_{t}-\widehat{\Delta}_{t}|$ (32)
where $y_{t}$ is the true value of the incidence at time $t$, and $\widehat{\Delta}_{t}=\mathrm{E}[\Delta_{t}|y_{1:t}]$ is the model posterior predictive mean of the incidence.
The CRPS evaluates the quality of the full predictive distribution by comparing the predicted cumulative distribution function (CDF) to the empirical CDF of the observation:
$\text{CRPS}=\frac{1}{T}\sum_{t=1}^{T}\int_{-\infty}^{\infty}\left(F_{t}(u)-\mathbb{I}(y_{t}\leq u)\right)^{2}\,du$
where $F_{t}(u)=p(\Delta_{t}\leq u\mid y_{1:T})$ is the predictive CDF at time $t$, and $\mathbb{I}(\cdot)$ is the indicator function.
When the predictive distribution is represented by a set of particles $\{x_{t}^{i}\}_{i=1}^{N_{x}}$, the CRPS can be approximated as:
$\text{CRPS}\approx\frac{1}{T}\sum_{t=1}^{T}\left(\frac{1}{N_{x}}\sum_{i=1}^{N_{x}}|y_{t}-x_{t}^{i}|-\frac{1}{2N_{x}^{2}}\sum_{i=1}^{N_{x}}\sum_{j=1}^{N_{x}}|x_{t}^{i}-x_{t}^{j}|\right)$
Figure 1: Daily estimates of the time-dependent transmission rate and effective reproduction number. The solid black lines represent the median estimate of $R_{\text{eff}}(t)$, with shaded blue regions indicating posterior uncertainty quantiles. Weekly averaged estimates of $R_{\text{eff}}(t)$ are shown in Figure 6.
Appendix E Application in real-time surveillance
The methodology developed in this paper is applied to a real-time surveillance setting, demonstrating how our model can track evolving dynamics in sequential data. By integrating new observations over time, the model provides insights into the system’s behavior. The focus is on an adaptive framework that updates the posterior distribution, conditioned on incoming data, while utilizing results from previous O-SMC2 runs as prior information. The results are shown in Figure 1, where we observe that there is no significant difference compared to the results obtained from running the model from scratch (see Figure 6).
Figure 1: Real-time application of O-SMC2 to sequential data: (Row 1) February 29, 2020, to May 30, 2020; (Row 2) May 30, 2020, to August 29, 2020; (Row 3) August 29, 2020, to November 28, 2020; and (Row 4) November 28, 2020, to February 27, 2021. Each row depicts the sequential estimates of daily new cases and the weekly average effective reproduction number ($R_{\text{eff}}(t)$). The solid black lines represent the median estimate, while the shaded blue regions indicate posterior uncertainty quantiles.